### v1

# EVE: Towards End-to-End Video Subtitle Extraction with Vision-Language Models

Abstract
The advent of Large Vision-Language Models (LVLMs) has advanced the video-based tasks, such as video captioning and video understanding. Some previous research indicates that taking texts in videos as input can further improve the performance of video understanding. As a type of indispensable information in short videos or movies, subtitles can assist LVLMs to better understand videos. Most existing methods for video subtitle extraction are based on a multi-stage framework, handling each frame independently. They can hardly exploit the temporal information of videos. Although some LVLMs exhibit the robust OCR capability, predicting accurate timestamps for subtitle texts is still challenging. In this paper, we propose an End-to-end Video Subtitle Extraction method, called EVE, which consists of three modules: a vision encoder, an adapter module, and a large language model. To effectively compress the visual tokens from the vision encoder, we propose a novel adapter InterleavedVT to interleave two modalities. It contains a visual compressor and a textual region compressor. The proposed InterleavedVT exploits both the merits of average pooling and Q-Former in token compression. Taking the temporal information of videos into account, we introduce a sliding-window mechanism in the textual region compressor. To benchmark the video subtitle extraction task, we propose a large dataset ViSa including 2.5M videos. Extensive experiments on ViSa demonstrate that the proposed EVE can outperform existing open-sourced tools and LVLMs.
1 Introduction
Due to the current prevalence of video as a mainstream media format, a significant amount of research is centered around video-related studies and developments, such as video understanding [34, 35, 37]. In the video understanding task, the text information present in videos, especially subtitle, plays an indispensable role since it provides more contextually relevant information of videos. Accurately extracting subtitles from videos can significantly assist existing Vision-Language Models (VLMs) in better understanding video content. Furthermore, video subtitle extraction can also provide foundational support for practical applications such as movie subtitle translation.
Figure 1: Most previous methods (a) rely on a multi-stage framework, where prediction error accumulates at each stage, influencing the final prediction. In contrast, the proposed EVE (b) is an end-to-end video subtitle extraction method.
Previous video subtitle extraction methods can be broadly categorized into two groups: OCR-based approach and speech-based approach. The OCR-based approach [7] often involves multiple stages, such as text detection, subtitle classification, and text recognition for subtitle regions (as shown in Fig. 1 (a)). Although these methods achieve reasonable performance, the multi-stage approach is susceptible to error accumulation. For example, if the output of text detection is inaccurate, text recognition models can hardly recognize corresponding texts correctly. In addition, they process each frame independently without considering the temporal information between frames. Different from the OCR-based approach, the speech-based approach [23, 22] directly transcribes the speech of videos to texts and timestamps the texts to generate subtitles. However, these methods are sensitive to noise in audios and often struggle to align the recognized subtitles with visual text cues in videos.
Recent advancements in VLMs have shown improved performance across various video tasks [12, 32, 38, 36]. Nevertheless, the field of video subtitle extraction still lacks an efficient end-to-end solution and corresponding datasets. In this paper, we introduce an End-to-end Video Subtitle Extraction method EVE, consisting of a vision encoder, an adapter module, and a Large Language Model (LLM). Specifically, we employ Swin-Transformer [18] as the vision encoder to extract visual features of each frame. Subsequently, we develop an innovative adapter module, named InterleavedVT, to compress the output visual tokens of the vision encoder. The proposed InterleavedVT consists of a visual compressor and a textual region compressor. The visual compressor is based on an average pooling operation to provide the global positional and visual information of subtitles within each video frame while the textual region compressor leverages a Q-Former [14] approach to extract subtitle-related tokens. Then, the output tokens of these two compressors are interleaved to construct a token sequence, which is further input into an LLM to predict subtitles. To support bilingual video subtitle extraction, we adopt Qwen2 [40] as the language model of our method. Additionally, to benchmark this task, we propose a large Video Subtitle Extraction dataset, called ViSa, based on existing two video datasets CNVid [9] and MovieChat [27]. The proposed ViSa includes 2.35M videos for pre-training, 230K videos for supervised fine-tuning, and 1.6K videos for test. The experimental results on ViSa demonstrate that EVE can achieve a significant performance improvement compared with previous open-sourced methods.
In summary, the contributions of this paper are as follows:
  • We first propose an end-to-end video subtitle extraction model EVE, consisting a vision encoder, an adapter module, and a large language model.
  • To condense the tokens of each frame, we develop a novel adapter InterleavedVT for video subtitle extraction. In InterleavedVT, a visual compressor and a textual region compressor are used to extract the global visual information and subtitle-related text information, respectively.
  • To benchmark the video subtitle extraction task, we propose a large video subtitle extraction dataset, called ViSa, including over 2.5M videos.
  • Compared with previous video subtitle extraction methods, the proposed EVE can achieve better performance in the test set of ViSa.
2 Related Work
2.1 Video Subtitle Extraction
Subtitles, as supplementary information in videos, play a crucial role in video understanding. Existing approaches to video subtitle extraction are mainly divided into two categories: OCR-based approach and speech-based approach. OCR-based methods typically treat video subtitle extraction as a multi-stage task. For instance,  [7] first perform text localization to locate the region of interest of frames. Then, it sequentially detects and recognizes subtitle texts in target frames.  [39] also adopts a similar pipeline, where Connectionist Text Proposal Network (CTPN) [28] is utilized for video subtitle detection. These methods consider information of each frame independently, struggling to leverage contextual information of videos. Additionally, errors in text detection and recognition will impact subtitle prediction. Speech-based methods [24] directly feed the audio content of videos into a speech recognition model, and further incorporate the transcribed texts with timestamps to generate corresponding subtitles. Since these methods treat the transcribed texts of the speech directly as subtitles, leading to discrepancies with the actual subtitles in videos. Especially in short videos (such as YouTube Shorts and TikTok videos), the alignment between subtitles and speech may not be perfect. Moreover, some open-source tools (e.g., VideOCR [^1] https://github.com/devmaxxing/videocr-PaddleOCR and PyVideoTrans [^2] https://github.com/jianchang512/pyvideotrans) or commercial cloud services (e.g., OpenVision [^3] https://vision.aliyun.com/ and GhostCut [^4] https://cn.jollytoday.com/home/) also support video subtitle extraction.
Figure 2: Overall architecture of EVE (a) and details of InterleavedVT (b). EVE contains three modules: a vision encoder, an adapter and a large language model. We propose a novel adapter called InterleavedVT, consisting of a visual compressor and a textual region compressor. To enhance the correlations between adjacent frames, we introduce a slide window mechanism.
2.2 Video Token Merging
Different from the tasks that take one single image as input, video-related tasks tend to involve more frames. Therefore, video-related Vision-Language Models (VLMs) need to employ token merging or compressing techniques to increase the number of input frames [13, 31, 45, 42]. Inspired by ToMe [5], many works have proposed various token merging methods to reduce the computational cost of models [16, 17]. Since video-related tasks often require leveraging the pre-trained vision encoder from image tasks, video-related tasks typically avoid compressing token quantities within the vision encoder and instead process the visual tokens output by the vision encoder. A common method is to employ a Q-Former-like module [14] to perform cross-attention operation between $N$ query tokens and visual tokens resulting $N$ compressed tokens, e.g., LLaMA-VID [15], TimeChat [26] and Video-LLaMA [43]. However, a recent research DeCo [41] indicates that only using Q-Former to compress visual tokens requires a large amount of data for training and may not converge well. In contrast, a simple average pooling may even perform better. Thus, some methods directly adopt an average pooling layer to compress input tokens, such as LITA [10], Video-ChatGPT [21] and Oryx MLLM [19]. To combine the metrits of both approaches, we propose Interleave VT, which achieves a high token compression rate while maintaining strong generalizability.
3 Methodology
3.1 Architecture
The overall architecture of the proposed End-to-end Video Subtitle Extraction method EVE is shown in Figure 2. Specifically, EVE consists of three modules: a vision encoder, an adapter and a large language model. Given an input video $\mathcal{V}=\{\mathbf{I}_{0},...,\mathbf{I}_{N-1}\}$, each frame $\mathbf{I}_{i}$ ($i\in\{0,1,...,N-1\}$) is independently input into the vision encoder to extract corresponding visual features $\mathbf{F}_{i}$. Thus, the output of the vision encoder can be denoted as =$\mathcal{F}=\{\mathbf{F}_{0},...,\mathbf{F}_{N-1}\}$, where $\mathbf{F}_{i}\in\mathbb{R}^{(H\times W)\times C}$ denotes the visual tokens of the $i^{th}$ frame. Subsequently, the visual features of all frames $\mathcal{F}$ are input into the proposed novel adapter InterleavedVT to produce the compressed visual tokens ^=$\hat{\mathcal{F}}=\{\hat{\mathbf{F}}_{0},...,\hat{\mathbf{F}}_{N-1}\}$. Finally, the visual tokens $\hat{\mathcal{F}}$ and embedded prompt $\mathcal{P}$ are fed into the large language model to generate corresponding subtitles $\mathcal{S}=\{S_{0},...,S_{T-1}\}$. The template of subtitle $S_{i}$ is “$\langle b_{i}\rangle\langle e_{i}\rangle s_{i}$”, where $b_{i}$/$e_{i}$ represents the begin/end index of frame for the $i$-th subtitle and $s_{i}$ is the corresponding text. In the following, we will detail each module in the proposed EVE.
Vision Encoder. In recent years, various Transformer-based vision backbones are proposed, including ViT [6], Swin-Transformer [18], EViT [16], Dynamic ViT [25], etc. Among these Transformer-based backbones, we select Swin-Transformer pretrained in Donut [11] as the vision encoder of EVE for two reasons: (1) Better initialization for OCR. Since the Swin-Transformer backbone of Donut has been pre-trained on numerous document datasets for multiple OCR tasks, it can provide better initialization to enhancing the text perception ability of EVE. In addition, based on the pre-training Swin-Transformer of Donut, our model can converge faster during the pre-training phase. (2) Higher time efficiency. Benefiting from the introduced shifted window-based self-attention module in Swin-Transformer, it shows better time efficiency, which can accelerate extracting visual features of frames in videos.
InterleavedVT. Although previous methods have proposed various adapter for video tasks, they may not suit for video subtitle task. To effectively compress the output visual tokens $\mathcal{F}$ in the video subtitle extraction task, we propose an innovative adapter InterleavedVT, which consists of a visual compressor and a textual region compressor. The details of InterleavedVT are shown in Figure 2. InterleavedVT fuses two types of token compression methods: average pooling and Q-Former. Due to the consistency of the subtitle positions in the video, the visual compressor is introduced to perform an average pooling operation on $\mathbf{F}_{i}$ to reserve the global visual features $\mathbf{F}^{v}_{i}\in\mathbb{R}^{4\times 4\times C}$ with relative positions. In addition, considering that recognizing texts of subtitles requires fine-grained textual information, we introduce the textual region compressor to reserve text-related tokens. Specifically, we randomly initialize $K$ queries $\mathbf{Q}$ to performance cross-attention with $\mathbf{F}_{i}$, resulting $K$ text-related tokens $\mathbf{F}^{t}_{i}\in\mathbb{R}^{K\times C}$, which can be defined as:
$\mathbf{F}^{t}_{i}=\texttt{CrossAttention}(Q,\mathbf{F}_{i})$ (1)
Taking into account the repeated subtitles between adjacent frames, we introduced a window mechanism in the textual region compressor to enhance the robustness of text-related token extraction. Thus, the processing of the textual region compressor can be modified into:
$\mathbf{F}^{t}_{i}=\texttt{CrossAttention}(Q,[\mathbf{F}_{i-1},\mathbf{F}_{i},\mathbf{F}_{i+1}])$ (2)
Similar to previous methods [26, 8], a projection layer is needed to align the features of vision encoder and language models. However, we observe that utilizing a single projection layer for both the visual and text-related tokens may lead to subpar performance (more discussions ). Since the two types of features from the visual compressor and the textual region compressor are regarded as different representations of one frame, we employ two projection layers to transform $\mathbf{F}^{v}_{i}$ and $\mathbf{F}^{t}_{i}$ into $\hat{\mathbf{F}}^{v}_{i}$ and $\hat{\mathbf{F}}^{t}_{i}$, respectively. Finally, the global visual tokens ^$\hat{\mathcal{F}}^{v}$ and the text-related tokens ^$\hat{\mathcal{F}}^{t}$ of all frames are interleaved into the output of InterleavedVT $\mathbf{F}^{out}$:
$\mathbf{F}^{out}=[\hat{\mathbf{F}}^{v}_{0},\hat{\mathbf{F}}^{t}_{0},...,\hat{\mathbf{F}}^{v}_{N-1},\hat{\mathbf{F}}^{t}_{N-1}]$ (3)
Input : Predicted subtitles $\mathcal{S}$={$S_{0}$, $S_{1}$, …, $S_{T-1}$}, where ; OCR results of all frame in the input video $\mathcal{O}$={$O_{0}$, $O_{1}$, …., $O_{L-1}$}, where $O_{l}$ = $\text{o}_{0}$, $\text{o}_{1}$,…, $\text{o}_{K_{l}-1}$}; Similarity threshold $\mathbf{sim}$, Raw FPS of the input video $\mathbf{f}$, Subtile matching range $\mathbf{r}$
Output : Subtitles in SRT format $\mathcal{S}^{\prime}$={$S^{\prime}_{0}$, $S^{\prime}_{1}$, …, $S^{\prime}_{T-1}$}, $S^{\prime}_{t}$=[$b^{\prime}_{t}$, $e^{\prime}_{t}$, $s^{\prime}_{t}$]
foreach subtitle S in $\mathcal{S}$ do
       # calculate the coarse timestamp of S ;
       $b_{c},e_{c}$ = $\frac{\mathbf{f}\times b}{\text{Sampling Rate}}$, $\frac{\mathbf{f}\times e}{\text{Sampling Rate}}$;
       # perform matching in the range $\mathbf{r}$ to locate the accurate start timestamp ;
       for $i$=$bc-\mathbf{r}$ to $b_{c}+\mathbf{r}$ do
             $min_{\text{ed}}$ = $\infty$;
             foreach Text o in $O_{i}$ do
[…]
                         $min_{\text{ed}}$ = EditDistance(s, o);
                   end if
             end foreach
            # locate the accurate start timestamp $b^{\prime}$ ;
             if $min_{\text{ed}}<\mathbf{sim}$ then
                   $b^{\prime}$ = i ;
                   break ;
             end if
       end for
      # perform the same operation as before to locate the accurate end timestamp $e^{\prime}$ ;
       # texts in predicted subtitles and converted subtitles are consistent ;
       $s^{\prime}$ = s
end foreach
return Subtitles in SRT format $\mathcal{S}^{\prime}$ ;
Algorithm 1 Convert EVE predictions into subtitles in SRT format.
Large Language Model. In the past few years, numerous foundation large language models are open-sourced, such as Qwen [40] and LLaMa [29]. Since the proposed EVE aims at solving bilingual (i.e., English and Chinese) video subtitle extraction, we select Qwen as the large language model in EVE. In addition, taking into account the time efficiency in practical applications, we build EVE-7B and EVE-0.5B based on Qwen2-7B and Qwen2-0.5B, respectively. Detail performance comparison is shown in Sec. 5.
3.2 Training
The training process of the proposed EVE is divided into two stages: pre-training stage and supervised fine-tuning stage. We compare these two different stages from the following three aspects:
  • Training Details. The pre-training stage primarily aims at enhancing the text perception capability of EVE, enabling it to predict all text tracklets in video frames. As shown in Figure 3, at the pre-training stage, the large language model is frozen while the vision encoder and the proposed adapter InterleavedVT are optimized. After the pre-training stage, the vision encoder is well trained to perceive texts. Thus, at the supervised fine-tuning stage, the vision encoder is frozen while the adapter InterleavedVT and the language model are optimized to learn discriminate whether texts in the input video are subtitles.
  • Input. Both pre-training and supervised fine-tuning stages share the same input template. In the template, we additionally prepend the corresponding frame index to the visual tokens to improve the temporal perception ability of our model. The limit for the number of input frames is set to 30 during pre-training while it is set to 120 during supervised fine-tuning.
  • Prediction. For prediction, the two stages also share the same template, where the start/end frame index and texts are included. Differently, the pre-training stage learns to recognize all text tracklets in videos while the supervised fine-tuning stage only considers those subtitle tracklets.
Figure 3: Training details of EVE. The LLM is frozen during pre-training while the vision encoder is forzen during supervised fine-tuning.
3.3 Inference
Due to the limitation of EVE in input token length, inference cannot be completed at high frame rates. Therefore, the proposed EVE samples the video at 2 frames per second (FPS) for model input. Although the predicted frame index can directly correspond to a precise timestamp (i.e., hour, minute, and second), it does not represent the precise time at which each subtitle appears and disappears in the video. To output more accurate subtitle content, such as in SRT format (examples are provided in the supplementary materials), we propose a post-processing method based on OCR results at full frame rate. The algorithm of post-processing is described in Algorithm 1. Firstly, we use an OCR engine to detect and recognize text in images at full frame rate. Then, we map the frame indices identified by EVE to rough timestamps at full frame rate. Finally, by calculating the edit distance between predicted texts and those within adjacent frames, the precise subtitle timestamp can be obtained.
It is worth mentioning that the aforementioned post-processing is not necessary for EVE. Due to the fact that taking videos with original frame rate as input will greatly decrease the video duration that EVE can handle, we extract frames of videos in 2-FPS as the input, which requires post-processing to output accurate timestamps. If we only input a two-second video at full frame rate (e.g., 120 frames in total), the model can generate accurate SRT subtitles without the need for post-processing. Therefore, it can be said that EVE is an end-to-end method for video subtitle extraction.
Chinese English Bilingual None-Text
Short Video 93,871 1,196 464 9,693
Movie 3,141 2,600 173 5,258
Table 1: The number of samples in different languages. Both short videos and movie videos are mainly in Chinese.
4 Data Construction
In this paper, we propose a large dataset, called ViSa, to benchmark the video subtitle extraction task. The raw videos are collect from two datasets: CNVid [9] and MovieChat-1K [27]. CNVid is the largest public video-text dataset in Chinese, which contains 3.5M short videos collect from Douyin  [^5] https://www.douyin.com/. MovieChat-1K contains 1K long movie videos with bilingual subtitles.
Based on these two datasets, we construct ViSa for video subtitle extraction. For CNVid, we perform the following three steps to reserve the ideal videos: (1) First, we filter out videos with a duration shorter than 10 seconds or longer than 120 seconds. The reason for this operation is that short videos (duration$<$10s) contain a lot of noise while the number of frames in long videos (duration$>$120s) exceeds the model’s input limits. (2) Then, we utilize a in-house video OCR tools to predict the text tracklets in the reserved videos. (3) Finally, videos with less than five tracklets will be filtered out, as these videos are unlikely to contain subtitles. Due to the fact that the movie videos in MovieChat-1K are mostly longer than 30 minutes, we clip each movie video into multiple short videos with a random duration (from 15s to 60s). By processing the above two datasets, we can obtain 2.5M short videos to build the video subtitle extraction dataset ViSa.
Figure 4: The statistics information of ViSa. The proportion of videos without subtitles in movie videos is higher that in short videos. Since we clip movie videos based on a uniformly sampled duration in the range of 15--60 seconds, the duration distribution of movie videos is more balanced compared to short videos.
Specifically, ViSa divides the collected videos into three parts: pre-training, supervised fine-tuning, and test datasets. Some statistics of them are described in the following. More details are shown in the supplementary materials.
Pre-training dataset. The pre-training dataset contains 2.35M videos, all of which are collected from CNVid. The videos in the pre-training dataset are annotated with text tracking and text recognition since the goal of pre-training is to equip the model with the text perception capability. The statistics of pre-training dataset are shown in the supplementary materials. The majority of pre-training videos are typically shorter than 60 seconds in length. The number of text tracklets in videos mostly ranges between 12 and 30. In a total of 2.5M pre-training videos, 1.8M videos include text tracklets while 0.7M videos contain no texts.
Supervised Fine-tuning (SFT) dataset. The videos for supervised fine-tuning originate from both CNVid and MovieChat-1K. Concretely, the SFT dataset includes 210K Chinese videos from CNVid and 20K bilingual videos from MovideChat-1K. Each video in the SFT dataset is annotated with subtitle text tracking and corresponding results. As a data augmentation strategy to expand the SFT dataset, we annotate each video with both 1-FPS and 2-FPS sampling rate. The statistics of SFT dataset are shown in Fig. 4. As shown in Tab. 1, among 230K SFT videos, 30K videos contains no texts.
Test dataset. To evaluate the performance of EVE on both short videos and movies, we collect 857 videos from CNVid and 862 videos from MovieChat-1K to build the test dataset. The annotations of this dataset are the same as that of the SFT dataset and they are manually annotated. Similar to the SFT dataset, both positive samples (i.e., videos containing texts) and negative samples (i.e., videos containing no texts) are also included in the test dataset.
Method Short Video Movie
Positive Negative Positive Negative
NED SubER SubER NED SubER SubER
Open-sourced Tools VideOCR [4] 0.348 72.72 52.63 0.686 6.28 57.12
VideOCR [4] + Subregion 0.659 52.92 43.81 0.702 5.24 66.35
PyVideoTrans [3] 0.477 46.03 92.34 0.639 88.35 52.51
Cloud Service OpenVision [2] 0.771 21.85 42.85 0.718 6.57 18.61
GhostCut [1] 0.655 54.78 46.00 0.659 7.26 24.43
Open-sourced LVLMs Qwen2-VL-2B [31] 0.068 - - 0.070 - -
Qwen2-VL-7B [31] 0.123 - - 0.227 - -
MiniCPM-V-2.6 [42] 0.043 - - 0.047 - -
Ours EVE-0.5B 0.759 22.82 18.36 0.685 5.10 17.06
EVE-7B 0.882 15.16 13.92 0.808 1.64 10.15
Table 2: Performance comparison with existing three types of methods. Since the open-sourced LVLMs are trained with no video subtitle extraction datasets, they cannot achieve ideal performance on this task.
5 Experiments
5.1 Implementation Details
The proposed method is implemented with PyTorch. All experiments are conducted on $24$ A100 GPUs. The proposed EVE is trained using the AdamW optimizer [20] where the learning rate is set to $5\times 10^{-5}$. The input frames are resized into $640\times 640$. The batch size is set to 48 for pre-training and 16 for supervised fine-tuning. The number of queries $K$ in the textual region compressor is set to 10, and the number of visual tokens is set to 16. More relevant experimental results are shown in the supplementary materials. The vision encoder is initialized with the pre-trained parameters of Donut [11], and the large language model loads the pre-trained models of Qwen2-7B or Qwen2-0.5B.
5.2 Metrics
In this paper, we adopt two types of metrics to evaluate the performance of models: Normalized Edit Distance (NED) [44] and SubER [33]. NED is defined as:
$\texttt{NED}(\hat{\mathbf{y}}_{i},\mathbf{y}_{i})=1-\frac{1}{S}\sum^{S}_{i=1}\texttt{ED}(\hat{\mathbf{y}}_{i},\mathbf{y}_{i})/\texttt{MaxL}(\hat{\mathbf{y}}_{i},\mathbf{y}_{i})$ (4)
where ED($\cdot$) denotes the edit distance function. The function MaxL(A, B) returns the maximum length between strings A and B. $S$ represents the number of samples. The NED metric is utilized to calculate the accuracy of predicting subtitle texts. However, it may hardly evaluate the accuracy of timestamp predictions. Therefore, we additionally adopt SubER [33] to evaluate the performance of models. SubER is defined as:
$\texttt{SUBER}=\frac{\#\text{word}\ \ \text{edits}+\#\text{break}\ \ \text{edits}+\#\text{shifts}}{\#\text{reference}\ \ \text{words}+\#\text{reference}\ \ \text{breaks}}$ (5)
where “#” means “number of”. word edits are insertions, deletions and substitutions of words; break edits are insertions, deletions and substitutions of line breaks; shifts are movements of one or more adjacent hypothesis tokens to a position of a matching phrase in the reference; reference words/reference breaks is the ground-truth words/breaks. More details about SubER can refer to [33].
5.3 Experimental Results
Evaluating Existing Methods. Although the video subtitle extraction task has been explored before, there is still a lack of comprehensive dataset to benchmark this task. In this paper, we propose a video subtitle extraction dataset containing pre-training, supervised fine-tuning and test datasets. We have evaluated three categories of video subtitle extraction methods on the proposed test dataset. Among the open-sourced tools, PyVideoTrans based on ASR achieves relatively inferior performance on negative samples compared with VideOCR. The reason for this results is that the audio in the negative sample will also be translated into sentences, which are regarded as subtitles of the input video. We have also evaluated two cloud services that can perform video subtitle extraction. Ghost cannot perform well on short videos containing a lot of background texts since it tends to mistake background texts as subtitles. In contrast, OpenVision can achieve relatively ideal performance on both short videos and movie videos. Two open-source multi-modal large models (Qwen2-VL [31] and MiniCPM-V-2.6 [42]) perform poorly on the video subtitle extraction task because they are not trained on corresponding datasets. They prefer to performing video captioning rather than extracting subtitle texts, thus resulting in subpar performance. In addition, the experimental results demonstrate that Qwen2-VL’s video OCR capability is better than MiniCPM-V-2.6. More visualizations are shown in supplementary materials.
Comparison with Existing Methods. In this paper, we establish two models with different LLMs. To achieve better performance in video subtitle extraction, we utilize Qwen2-7B to construct EVE-7B. Through the experimental results in Tab. 2, EVE-7B achieves the best performance compared with three types of video subtitle extraction methods. For example, on the short video dataset, EVE-7B outperforms OpenVision by 11.1% in the metric of NED. We have visualized an example in Fig 5, and the result shows that the proposed EVE can well predict the subtitles without missing and repetition problem. More visualizations are shown in the supplementary materials. We have also conducted experiments to compare with two typical token compression methods (T-Selector [30] and ToMe [5]). The results are shown in the supplementary materials.
Considering that the EVE-7B will cost more computing resources in practical applications, we establish a light-weighted model EVE-0.5B with Qwen2-0.5B. The experimental results in Tab. 2 demonstrate that despite using the same dataset for training, EVE-0.5B still cannot achieve the same performance as EVE-7B. However, benefiting from the proposed InterleavedVT and corresponding dataset ViSa, it can still surpass most of existing methods.
Figure 5: An example of subtitle extraction for different methods. The ASR based method PyVideoTrans tends to miss some audio-irrelative texts; Qwen2-VL performs video captioning rather than extracting subtitles; OpenVision may be based on text tracking, resulting repetition problem when text tracklets are wrongly predicted. Gray and Red indicate the missing and repetition prediction problem, respectively. Green represents the format of the answer is incorrect.
VC TRC SP Short Video Movie
Pos Neg Pos Neg
18.15 19.34 7.25 14.26
22.35 20.19 9.41 17.80
17.32 16.28 4.17 15.82
15.16 13.92 1.64 10.15
Table 3: Ablation studies on the proposed EVE (metric: SubER). Through experimental results, the separate projector is necessary for InterleavedVT.
5.4 Ablation Studies
In this paper, we propose an innovative adapter called InterleavedVT, which includes a visual compressor (VC) and a textual region compressor (TRC) for compressing visual features. Additionally, we introduce the separate projector (SP) to project the visual tokens and text-related tokens into the input space of large language models. In this section, we conduct ablation studies on these three modules. The experimental results, as shown in Tab. 3, indicate that introducing either the visual compressor or the textual region compressor alone does not outperform previous methods. However, when both of them are integrated into InterleavedVT, there is a significant improvement in model performance. Compared to only using the textual region compressor, our method achieves a 5.03% and 5.28% improvement in SubER metrics on the positive samples of short video and movie datasets, respectively. Furthermore, introducing separate projectors can further improve the performance. A plausible explanation is that the two types of information focus on different aspects of visual effects, and thus, separate projectors facilitate better alignment of features with LLMs.
6 Limitations and Future Work
Although the proposed EVE have compressed each frame into 26 tokens and supported to input 120 frames for video subtitle extraction, it still have difficulties in solving long videos (e.g., longer than one hour). In future work, we will explore to more effectively compress text-related tokens in video frames to increase the number of image inputs to the model. Furthermore, the proposed EVE can only predict timestamps and texts of subtitles, ignoring the positional information of texts (i.e., bounding box). To support more downstream applications, we will attempt to predict the accurate coordinates of subtitles in each frame.
7 Conclusion
In this paper, we propose an End-to-end Video Subtitle Extraction model, called EVE, which consists of a vision encoder, an adapater module and a large language model. In EVE, a novel adapter, called InterleavedVT, is introduced for effectively compressing visual tokens. To benchmark the video subtitle extraction task, we propose a large bilingual (i.e., Chinese and English) dataset ViSa, including 2.35M videos for pre-training, 230K videos for fine-tuning and 1.6K videos for test. Through experimental results, we observe that the proposed EVE achieves the best performance in both NED and SubER compared with previous methods for video subtitle extraction.
Supplementary Material
8 SRT Format
An SRT file is one of the most popular subtitle file formats for video content. These plain text files include the text of the subtitles in sequence, along with the start and end timestamps. However, an SRT file does not contain any video or audio content. That’s why the file sizes are so small. An SRT file needs to specify the line number, subtitle content, when the text should appear on the video, and when it should disappear to make room for the next line. An example is shown in the following.
Figure 6: An example of SRT files.
9 Hyperparameter Selection
There are two hyperparameters in EVE: the number of queries in the textual region compressor and the number of visual tokens. We have conducted corresponding experiments with different numbers of tokens. When the number of visual tokens (denoted as $N$) and the number of queries in the textual region compressor (denoted as $K$) are set to larger values of ($N$=36 and $K$=20), the performance of EVE is only improved by 0.3% in the NED metric, but the computational cost will increase fourfold. When $N$ and $K$ are set to smaller values ($N$=4 and $K$=1), there is a significant decrease in the performance of EVE, with the NED metric decreasing by nearly 28.6%. In order to balance computational cost and model performance, we ultimately set $N$ and $K$ to 16 and 10, respectively.
10 Comparison with Token Merging Methods
Except the open-sourced subtitle extraction tools and LVLMs, we also compare our method with two typical token Merging methods (T-Selector [30] and ToMe [5]). The results are shown in Tab. 4. We use Qwen2-7B to implement T-Selector and ToMe for subtitle extraction. The experimental results show that although our method EVE-0.5B only adopts a 0.5B language model, it can achieve comparable results to T-Selector and ToMe. When we use Qwen2-7B for EVE, our method can achieve significant performance improvements. For example, compared to ToMe , EVE can achieve a 16.8% performance improvement (in NED) on the Short Videos test dataset. The results demonstrate that the proposed token merging method InterleavedVT can more efficiently extract the subtitle-related information in videos.
Method Short Video Movie
Positive Negative Positive Negative
NED SubER SubER NED SubER SubER
T-Selector [30] 0.801 20.32 23.84 0.731 8.72 19.30
ToMe [5] 0.714 28.35 25.62 0.705 10.79 15.59
EVE-0.5B 0.759 22.82 18.36 0.685 5.10 17.06
EVE-7B 0.882 15.16 13.92 0.808 1.64 10.15
Table 4: Performance comparison with two typical token merging methods.
11 Visualization for Textual Region Compressor
In the proposed EVE, the textual region compressor is introduced to perceive text-related information in each frame. To verify the effectiveness of this module, we have visualized the attention map of the textual region compressor in Fig. 7. In the visualization, we observe that this module can indeed focus on the text region in each frame, which demonstrates the effectiveness of the proposed textual region compressor.
Figure 7: The visualization for the textual region compressor.
12 More Statistical Information of ViSa
The detail information of SFT dataset have shown in the main text. In this section, we show some statistical information of the pre-training data in Fig. 8. Most of pre-training videos contain less than 10 subtitles and their duration is between 11-20 seconds.
13 More Result Visualizations
We additionally visualize two comparison examples. Fig.9 and Fig.10 show the short video and movie examples, respectively. Through these two examples, we observe that Qwen2-VL and MiniCPM tend to describe the video content rather than predicting subtitles in SRT format. In addition, the cloud services can hardly extract bilingual subtitles, especially in movie videos. The ASR-based method PyVideoTrans also cannot achieve ideal performance since it only takes audio as input, which is not fully aligned with visual information. Compared with previous methods, the proposed EVE can better predict texts and timestamps of subtitles.
Figure 8: The statistics information of ViSa’s pre-training dataset.
Figure 9: A comparison example of short video.
Figure 10: A comparison example of movie.

### v2

@@
 Abstract
-The advent of Large Vision-Language Models (LVLMs) has advanced the video-based tasks, such as video captioning and video understanding.
-Some previous research indicates that taking texts in videos as input can further improve the performance of video understanding.
-As a type of indispensable information in short videos or movies, subtitles can assist LVLMs to better understand videos.
-Most existing methods for video subtitle extraction are based on a multi-stage framework, handling each frame independently.
-They can hardly exploit the temporal information of videos.
-Although some LVLMs exhibit the robust OCR capability, predicting accurate timestamps for subtitle texts is still challenging.
-In this paper, we propose an End-to-end Video Subtitle Extraction method, called EVE, which consists of three modules: a vision encoder, an adapter module, and a large language model.
-To effectively compress the visual tokens from the vision encoder, we propose a novel adapter InterleavedVT to interleave two modalities.
-It contains a visual compressor and a textual region compressor.
-The proposed InterleavedVT exploits both the merits of average pooling and Q-Former in token compression.
-Taking the temporal information of videos into account, we introduce a sliding-window mechanism in the textual region compressor.
-To benchmark the video subtitle extraction task, we propose a large dataset ViSa including 2.5M videos.
-Extensive experiments on ViSa demonstrate that the proposed EVE can outperform existing open-sourced tools and LVLMs.
-1 Introduction
-Due to the current prevalence of video as a mainstream media format, a significant amount of research is centered around video-related studies and developments, such as video understanding [34, 35, 37].
-In the video understanding task, the text information present in videos, especially subtitle, plays an indispensable role since it provides more contextually relevant information of videos.
-Accurately extracting subtitles from videos can significantly assist existing Vision-Language Models (VLMs) in better understanding video content.
-Furthermore, video subtitle extraction can also provide foundational support for practical applications such as movie subtitle translation.
-Figure 1: Most previous methods (a) rely on a multi-stage framework, where prediction error accumulates at each stage, influencing the final prediction.
+Video subtitles play a crucial role in short videos and movies, as they not only help models better understand video content but also support applications such as video translation and content retrieval.
+Existing video subtitle extraction methods typically rely on multi-stage frameworks, where errors accumulate across stages and temporal dependencies are underutilized due to frame-wise processing.
+Moreover, although some Large Vision-Language Models (LVLMs) possess strong OCR capabilities, predicting accurate timestamps for subtitle texts remains challenging.
+To this end, we propose an End-to-end Video subtitle Extraction framework based on LVLMs, named EVE, which can output subtitles and their timestamps simultaneously.
+Specifically, we introduce a dual-branch Spatiotemporal Subtitle-Salient (S3) Module that serves as an adapter for LVLMs, capable of representing subtitle-related content and considering inter-frame correlations using only a small number of tokens.
+Within this module, the Spatial Semantic Context Aggregate branch aggregates high-level global semantics to provide spatial visual contextual information, while the Temporal Subtitle Token Query branch explicitly queries subtitle-relevant tokens while considering temporal correlation across frames.
+The small number of tokens retained by the S3 module are fed to the language model, which then directly outputs the subtitle text along with its timestamps.
+Furthermore, we construct the first large-scale dataset dedicated to video subtitle extraction, ViSa, containing over 2.5M videos with timestamped and bilingual annotation, thereby providing the community with a well-organized training and evaluation benchmark.
+Extensive experiments on ViSa demonstrate that EVE significantly outperforms existing open-source tools and LVLMs in both subtitle recognition accuracy and temporal alignment, achieving state-of-the-art performance.
+I Introduction
+With the rapid growth of video as the dominant medium for information dissemination, video understanding has become a central topic in multi-modal research [wu2021towards, wu2024deep, tip-video-sum1, tip-video-sum2].
+Among the various modalities embedded in videos, textual information—particularly subtitles—serves as an indispensable bridge between visual and linguistic domains, providing explicit and contextually aligned semantic information.
+Accurate subtitle extraction not only enhances the ability of Vision-Language Models (VLMs) to comprehend video content but also enables a wide range of downstream applications, such as cross-lingual subtitle translation and video content retrieval [tip-video-ret].
+Therefore, video subtitle extraction is of great importance both as a fundamental research problem and as a key enabling technology for practical video understanding applications.
+Figure 1: Comparison of the proposed method with previous methods on the pipeline.
+Most previous methods (a) rely on a multi-stage framework, where prediction error accumulates at each stage, influencing the final prediction.
 In contrast, the proposed EVE (b) is an end-to-end video subtitle extraction method.
 Previous video subtitle extraction methods can be broadly categorized into two groups: OCR-based approach and speech-based approach.
-The OCR-based approach [7] often involves multiple stages, such as text detection, subtitle classification, and text recognition for subtitle regions (as shown in Fig.
-1 (a)).
+The OCR-based approach [elshahaby2022end, yan2020end] often involves multiple stages, such as text detection, temporal tracking, and text recognition for subtitle regions as shown in Figure 1 (a).
 Although these methods achieve reasonable performance, the multi-stage approach is susceptible to error accumulation.
-For example, if the output of text detection is inaccurate, text recognition models can hardly recognize corresponding texts correctly.
+For example, if the output of text detection is inaccurate, text recognition models can hardly recognize the corresponding texts correctly.
 In addition, they process each frame independently without considering the temporal information between frames.
-Different from the OCR-based approach, the speech-based approach [23, 22] directly transcribes the speech of videos to texts and timestamps the texts to generate subtitles.
+Different from the OCR-based approach, the speech-based approach [papi2023direct, mathur2015generating] directly transcribes the speech of videos to texts and timestamps the texts to generate subtitles.
 However, these methods are sensitive to noise in audios and often struggle to align the recognized subtitles with visual text cues in videos.
-Recent advancements in VLMs have shown improved performance across various video tasks [12, 32, 38, 36].
-Nevertheless, the field of video subtitle extraction still lacks an efficient end-to-end solution and corresponding datasets.
-In this paper, we introduce an End-to-end Video Subtitle Extraction method EVE, consisting of a vision encoder, an adapter module, and a Large Language Model (LLM).
-Specifically, we employ Swin-Transformer [18] as the vision encoder to extract visual features of each frame.
-Subsequently, we develop an innovative adapter module, named InterleavedVT, to compress the output visual tokens of the vision encoder.
-The proposed InterleavedVT consists of a visual compressor and a textual region compressor.
-The visual compressor is based on an average pooling operation to provide the global positional and visual information of subtitles within each video frame while the textual region compressor leverages a Q-Former [14] approach to extract subtitle-related tokens.
-Then, the output tokens of these two compressors are interleaved to construct a token sequence, which is further input into an LLM to predict subtitles.
-To support bilingual video subtitle extraction, we adopt Qwen2 [40] as the language model of our method.
-Additionally, to benchmark this task, we propose a large Video Subtitle Extraction dataset, called ViSa, based on existing two video datasets CNVid [9] and MovieChat [27].
-The proposed ViSa includes 2.35M videos for pre-training, 230K videos for supervised fine-tuning, and 1.6K videos for test.
-The experimental results on ViSa demonstrate that EVE can achieve a significant performance improvement compared with previous open-sourced methods.
-In summary, the contributions of this paper are as follows:
-  • We first propose an end-to-end video subtitle extraction model EVE, consisting a vision encoder, an adapter module, and a large language model.
-  • To condense the tokens of each frame, we develop a novel adapter InterleavedVT for video subtitle extraction.
-In InterleavedVT, a visual compressor and a textual region compressor are used to extract the global visual information and subtitle-related text information, respectively.
-  • To benchmark the video subtitle extraction task, we propose a large video subtitle extraction dataset, called ViSa, including over 2.5M videos.
-  • Compared with previous video subtitle extraction methods, the proposed EVE can achieve better performance in the test set of ViSa.
-2 Related Work
-2.1 Video Subtitle Extraction
+Recent advances in Vision-Language Models (VLMs) have substantially improved the performance of various video-related tasks [kim2024image, wang2024videotree, xu2024pllava, wu2023bidirectional].
+However, the task of video subtitle extraction remains underexplored.
+To our knowledge, there is no efficient end-to-end VLM-based frameworks or comprehensive datasets available.
+In addition, video subtitle extraction poses two unique challenges that distinguish it from general video understanding.
+First, subtitles appear in very small spatial regions and depend on fine-grained visual cues, which are easily lost under the token compression and global pooling strategies typical of existing LVLMs.
+Second, subtitles persist across multiple frames with gradual transitions, requiring high-fidelity modeling of both their textual content and temporal boundaries.
+Such temporal coherence cannot be captured by frame-wise OCR or captioning, often resulting in inconsistent or misaligned subtitles.
+To fill this gap, we propose an End-to-end Video subtitle Extraction framework, termed EVE.
+EVE comprises three major components: a vision encoder, a Spatiotemporal Subtitle-Salient (S3) module, and a large language model (LLM).
+The S3 module consists of two branch and acts as an adapter that refines frame-level visual features from two complementary perspectives while retaining only a small number of tokens.
+Specifically, the Spatial Semantic Context Aggregate (SSCA) branch aggregate high-level global semantics to provide spatial contextual cues for subtitle extraction.
+Meanwhile, the Temporal Subtitle Token Query (TSTQ) branch explicitly queries subtitle-relevant tokens inspired by the Q-former [li2023blip].
+Q-former is designed for general vision-language alignment, while TSTQ explicitly targets subtitle-bearing regions and enforces temporal consistency.
+The outputs from both branches are individually projected into the language model’s embedding space and then concatenated before being fed into the LLM to generate timestamped subtitle texts.
+To support bilingual (i.e., Chinese and English) video subtitle extraction, we employ Qwen2 [yang2024qwen2] as the large language model, leveraging its strong cross-lingual understanding and generation capabilities to ensure subtitle accuracy.
+Although several datasets [elshahaby2022end, 1227749, 6628859, Wang20111457] have been used in existing video subtitle extraction method, they suffer from issues such as limited scale, lack of timestamp annotations, or derivation from images rather than videos.
+This renders these datasets vastly inferior to those for other relevant tasks, rendering them inadequate to meet the current data requirements of LVLM.
+Therefore, we construct a large-scale benchmark dataset, ViSa, containing over 2.5M videos collected from CNVid [gan2023cnvid] and MovieChat [song2024moviechat] for pre-training, supervised fine-tuning, and testing.
+ViSa provides bilingual annotations of subtitle texts and timestamps, offering a valuable resource for systematic evaluation and future research on video subtitle extraction.
+Extensive experiments on ViSa demonstrate that EVE significantly outperforms existing open-source tools and LVLM baselines in both subtitle recognition accuracy and temporal alignment, setting a new state-of-the-art for video subtitle extraction.
+In summary, the main contributions of this work are as follows:
+  • We propose an End-to-end Video subtitle Extraction framework, EVE, built upon LVLMs to directly generate timestamped subtitles from video input.
+  • We design a Spatiotemporal Subtitle-Salient (S3) Module, which refines visual features through spatial semantic aggregation and temporal subtitle token querying to effectively capture subtitle-salient representations and consider temporal consistency with only a small number of tokens.
+  • We construct the first large-scale dataset, ViSa, containing over 2.5M videos with bilingual subtitle text and timestamp annotations, providing a valuable resource for systematic evaluation and future research in this field.
+  • Extensive experiments on ViSa demonstrate that EVE significantly outperforms existing open-source tools and LVLM baselines in video subtitle extraction task.
+The remainder of this paper is organized as follows: Section II reviews existing work on video subtitle extraction.
+Section III presents the overall architecture of EVE, including the vision encoder, the proposed S3 module, and the training and inference procedures.
+Section IV outlines the characteristics of the ViSa dataset and its construction process.
+Section V reports extensive experiments, including comparisons with existing tools and LVLMs, ablation studies, visualizations, and hyperparameter analyses.
+Finally, Section VI concludes the paper, discusses the limitations of the proposed method, and outlines potential future directions.
+II Related Work
+II-A Video Subtitle Extraction
 Subtitles, as supplementary information in videos, play a crucial role in video understanding.
-Existing approaches to video subtitle extraction are mainly divided into two categories: OCR-based approach and speech-based approach.
-OCR-based methods typically treat video subtitle extraction as a multi-stage task.
-For instance,  [7] first perform text localization to locate the region of interest of frames.
-Then, it sequentially detects and recognizes subtitle texts in target frames.
-[39] also adopts a similar pipeline, where Connectionist Text Proposal Network (CTPN) [28] is utilized for video subtitle detection.
-These methods consider information of each frame independently, struggling to leverage contextual information of videos.
-Additionally, errors in text detection and recognition will impact subtitle prediction.
-Speech-based methods [24] directly feed the audio content of videos into a speech recognition model, and further incorporate the transcribed texts with timestamps to generate corresponding subtitles.
-Since these methods treat the transcribed texts of the speech directly as subtitles, leading to discrepancies with the actual subtitles in videos.
-Especially in short videos (such as YouTube Shorts and TikTok videos), the alignment between subtitles and speech may not be perfect.
-Moreover, some open-source tools (e.g., VideOCR [^1] https://github.com/devmaxxing/videocr-PaddleOCR and PyVideoTrans [^2] https://github.com/jianchang512/pyvideotrans) or commercial cloud services (e.g., OpenVision [^3] https://vision.aliyun.com/ and GhostCut [^4] https://cn.jollytoday.com/home/) also support video subtitle extraction.
-Figure 2: Overall architecture of EVE (a) and details of InterleavedVT (b).
+Existing approaches to video subtitle extraction can be broadly categorized into two groups: OCR-based and speech-based approaches.
+II-A1 OCR-based approach
+OCR-based approaches typically treat subtitle extraction as a multi-stage pipeline.
+For example, Elshahaby et al.
+[elshahaby2022end] proposed an system that first performs text localization to identify regions of interest within video frames, followed by text detection and recognition on the localized areas.
+Similarly, Yan et al.
+[yan2020end] adopted a comparable framework, employing the Connectionist Text Proposal Network (CTPN) [tian2016detecting] for subtitle region detection.
+However, such methods process each frame independently, making it difficult to leverage the temporal and contextual information across frames.
+Moreover, the sequential nature of detection and recognition stages often leads to error accumulation, where inaccurate detection can significantly degrade the final recognition performance.
+VideOCR [^1] https://github.com/devmaxxing/videocr-PaddleOCR is an open-source video subtitle extraction tool developed based on the OCR tool PaddleOCR  [^2] https://github.com/PaddlePaddle/PaddleOCR.
+II-A2 Speech-based approach
+Mathur et al.
+[mathur2015generating] presents an offline pipeline that extracts audio from video, applies speech recognition, and converts the recognized speech into time-aligned subtitle files.
+Ramani et al.
+[ramani2020automatic] extract subtitles by transcribing the audio stream of videos using speech recognition models and then associating the transcribed text with timestamps.
+Sara et al.
+[papi2023direct] proposes an end-to-end direct speech translation model that generates time-aligned target-language subtitles from source audio.
+While effective for clear speech, these methods directly treat transcribed speech as subtitles, which can cause mismatches with the actual on-screen text—particularly in short-form videos (e.g., YouTube Shorts, TikTok) where subtitle alignment with speech is often imperfect.
+In addition to the aforementioned two types of methods, there are also tools available for extracting video subtitles.
+PyVideoTran  [^3] https://github.com/jianchang512/pyvideotrans is a fully automatic video/audio translation tool that transcribes speech, generates and translates subtitles, which also supports batch processing of video or audio files into precisely time-coded SRT subtitle files.
+Commercial cloud services (e.g., OpenVision [^4] https://vision.aliyun.com/, GhostCut [^5] https://cn.jollytoday.com/home/) have also been developed for video subtitle extraction.
+II-B Large Vision-Language Models
+In recent years, Large Vision-Language Models (LVLMs) [wang2024qwen2, yao2024minicpm, zhang2023video, li2025llama, huang2025lita] have demonstrated remarkable capabilities in instruction following and multimodal reasoning, enabling them to capture on-screen text or subtitles when prompted with appropriate instructions.
+Unlike conventional OCR-based methods, these models can leverage global visual-semantic context and learn high-level associations between video content and textual cues.
+However, their subtitle extraction performance remains limited in both accuracy and temporal alignment, mainly due to the lack of task-specific training for fine-grained subtitle localization.
+Most existing LVLMs are optimized for holistic video captioning or question answering [song2024moviechat, wang2024qwen2] rather than precise frame-level text prediction.
+Consequently, they tend to produce scene-level descriptions or paraphrased summaries instead of exact subtitle transcripts.
+This highlights the need for a task-specific framework that can jointly model spatial and temporal semantics to achieve accurate and temporally consistent subtitle extraction.
+Figure 2: Overall architecture of EVE (a) and details of S3 module (b).
 EVE contains three modules: a vision encoder, an adapter and a large language model.
-We propose a novel adapter called InterleavedVT, consisting of a visual compressor and a textual region compressor.
-To enhance the correlations between adjacent frames, we introduce a slide window mechanism.
-2.2 Video Token Merging
-Different from the tasks that take one single image as input, video-related tasks tend to involve more frames.
-Therefore, video-related Vision-Language Models (VLMs) need to employ token merging or compressing techniques to increase the number of input frames [13, 31, 45, 42].
-Inspired by ToMe [5], many works have proposed various token merging methods to reduce the computational cost of models [16, 17].
-Since video-related tasks often require leveraging the pre-trained vision encoder from image tasks, video-related tasks typically avoid compressing token quantities within the vision encoder and instead process the visual tokens output by the vision encoder.
-A common method is to employ a Q-Former-like module [14] to perform cross-attention operation between $N$ query tokens and visual tokens resulting $N$ compressed tokens, e.g., LLaMA-VID [15], TimeChat [26] and Video-LLaMA [43].
-However, a recent research DeCo [41] indicates that only using Q-Former to compress visual tokens requires a large amount of data for training and may not converge well.
-In contrast, a simple average pooling may even perform better.
-Thus, some methods directly adopt an average pooling layer to compress input tokens, such as LITA [10], Video-ChatGPT [21] and Oryx MLLM [19].
-To combine the metrits of both approaches, we propose Interleave VT, which achieves a high token compression rate while maintaining strong generalizability.
-3 Methodology
-3.1 Architecture
-The overall architecture of the proposed End-to-end Video Subtitle Extraction method EVE is shown in Figure 2.
-Specifically, EVE consists of three modules: a vision encoder, an adapter and a large language model.
-Given an input video $\mathcal{V}=\{\mathbf{I}_{0},...,\mathbf{I}_{N-1}\}$, each frame $\mathbf{I}_{i}$ ($i\in\{0,1,...,N-1\}$) is independently input into the vision encoder to extract corresponding visual features $\mathbf{F}_{i}$.
-Thus, the output of the vision encoder can be denoted as =$\mathcal{F}=\{\mathbf{F}_{0},...,\mathbf{F}_{N-1}\}$, where $\mathbf{F}_{i}\in\mathbb{R}^{(H\times W)\times C}$ denotes the visual tokens of the $i^{th}$ frame.
-Subsequently, the visual features of all frames $\mathcal{F}$ are input into the proposed novel adapter InterleavedVT to produce the compressed visual tokens ^=$\hat{\mathcal{F}}=\{\hat{\mathbf{F}}_{0},...,\hat{\mathbf{F}}_{N-1}\}$.
-Finally, the visual tokens $\hat{\mathcal{F}}$ and embedded prompt $\mathcal{P}$ are fed into the large language model to generate corresponding subtitles $\mathcal{S}=\{S_{0},...,S_{T-1}\}$.
-The template of subtitle $S_{i}$ is “$\langle b_{i}\rangle\langle e_{i}\rangle s_{i}$”, where $b_{i}$/$e_{i}$ represents the begin/end index of frame for the $i$-th subtitle and $s_{i}$ is the corresponding text.
-In the following, we will detail each module in the proposed EVE.
-Vision Encoder.
-In recent years, various Transformer-based vision backbones are proposed, including ViT [6], Swin-Transformer [18], EViT [16], Dynamic ViT [25], etc.
-Among these Transformer-based backbones, we select Swin-Transformer pretrained in Donut [11] as the vision encoder of EVE for two reasons: (1) Better initialization for OCR.
+We propose a novel adapter called S3 module, consisting of a spatial semantic context aggregating and a temporal subtitle token querying branches.
+To enhance the correlations between adjacent frames, we also introduce a temporal window mechanism.
+II-C Video Token Compression
+The proposed S3 module aims to extract subtitle-relevant representations from dense visual tokens and retain only a compact set of key tokens for the large language model.
+This mechanism is conceptually related to the token compression strategies widely adopted in recent LVLMs, which also focus on preserving the most informative tokens while reducing redundancy to enhance efficiency.
+Existing token compression methods can be roughly divided into two categories.
+The first category relies on token merging, where similar visual tokens are grouped or averaged to reduce redundancy, as exemplified by ToMe [bolyatoken] and its variants [liang2022not, liusimple].
+The second category employs query-based token selection, which introduces a small set of learnable queries to extract compact visual representations through cross-attention, similar to the Q-Former module [li2023blip].
+This approach has been adopted by several video-related VLMs such as LLaMA-VID [li2025llama], TimeChat [ren2024timechat], and Video-LLaMA [zhang2023video].
+However, studies such as DeCo [yao2024deco] have shown that Q-Former-based compression often requires large-scale training data and can suffer from unstable convergence.
+In contrast, simple operations like average pooling have demonstrated surprisingly strong performance in some cases, as reported in LITA [huang2025lita], Video-ChatGPT [maaz2023video], and Oryx  [liu2024oryx].
+Although our S3 module shares the same philosophy of extracting a minimal yet informative token set, it is specifically designed for the video subtitle extraction task.
+Instead of general-purpose visual redundancy reduction, S3 focuses on identifying tokens that are semantically and temporally relevant to subtitles.
+III Methodology
+III-A Architecture
+The overall architecture of the proposed End-to-end Video subtitle Extraction framework (EVE) is illustrated in Figure 2.
+EVE consists of three main components: a vision encoder, a Spatiotemporal Subtitle-Salient (S3) Module, and a large language model (LLM).
+Given an input video $\mathcal{V}=\{\mathbf{I}_{0},...,\mathbf{I}_{N-1}\}$, each frame $\mathbf{I}_{i}$ ($i\in\{0,1,...,N-1\}$) is first processed by the vision encoder to extract frame-level visual features $\mathbf{F}_{i}$.
+The resulting feature sequence can be denoted as $\mathcal{F}=\{\mathbf{F}_{0},...,\mathbf{F}_{N-1}\}$, where $\mathbf{F}_{i}\in\mathbb{R}^{(H\times W)\times C}$ represents the dense visual tokens of the $i$-th frame.
+To obtain subtitle-relevant representations, we propose the S3 module, which comprises two components: the Spatial Semantic Context Aggregation (SSCA) and the Temporal Subtitle Token Query (TSTQ).
+The former aggregates global semantics from spatial contexts, while the latter queries subtitle-relevant tokens through temporal salience modeling.
+The output of S3 is projected into the language model’s embedding space and denoted as $\mathbf{F}^{out}$ after concatenation.
+This representation, conditioned on an instruction prompt $\mathcal{P}$, is subsequently fed into the large language model to produce the subtitle sequence $\mathcal{S}={S_{0},\ldots,S_{T-1}}$, where each subtitle $S_{i}$ is formatted as $\langle b_{i}\rangle\langle e_{i}\rangle s_{i}$”, with $b_{i}$ and $e_{i}$ denoting its start and end frames.
+In the following subsections, we describe the vision encoder, the proposed S3 Module, and the large language model in detail.
+III-B Vision Encoder
+In recent years, various Transformer-based vision backbones are proposed, including ViT [dosovitskiy2020image], Swin-Transformer [liu2021swin], EViT [liang2022not], Dynamic ViT [rao2021dynamicvit], etc.
+Among these Transformer-based backbones, we select Swin-Transformer pretrained in Donut [kim2021donut] as the vision encoder of EVE for two reasons: (1) Better initialization for OCR.
 Since the Swin-Transformer backbone of Donut has been pre-trained on numerous document datasets for multiple OCR tasks, it can provide better initialization to enhancing the text perception ability of EVE.
@@
 Benefiting from the introduced shifted window-based self-attention module in Swin-Transformer, it shows better time efficiency, which can accelerate extracting visual features of frames in videos.
-InterleavedVT.
-Although previous methods have proposed various adapter for video tasks, they may not suit for video subtitle task.
-To effectively compress the output visual tokens $\mathcal{F}$ in the video subtitle extraction task, we propose an innovative adapter InterleavedVT, which consists of a visual compressor and a textual region compressor.
-The details of InterleavedVT are shown in Figure 2.
-InterleavedVT fuses two types of token compression methods: average pooling and Q-Former.
-Due to the consistency of the subtitle positions in the video, the visual compressor is introduced to perform an average pooling operation on $\mathbf{F}_{i}$ to reserve the global visual features $\mathbf{F}^{v}_{i}\in\mathbb{R}^{4\times 4\times C}$ with relative positions.
-In addition, considering that recognizing texts of subtitles requires fine-grained textual information, we introduce the textual region compressor to reserve text-related tokens.
-Specifically, we randomly initialize $K$ queries $\mathbf{Q}$ to performance cross-attention with $\mathbf{F}_{i}$, resulting $K$ text-related tokens $\mathbf{F}^{t}_{i}\in\mathbb{R}^{K\times C}$, which can be defined as:
-$\mathbf{F}^{t}_{i}=\texttt{CrossAttention}(Q,\mathbf{F}_{i})$ (1)
-Taking into account the repeated subtitles between adjacent frames, we introduced a window mechanism in the textual region compressor to enhance the robustness of text-related token extraction.
-Thus, the processing of the textual region compressor can be modified into:
-$\mathbf{F}^{t}_{i}=\texttt{CrossAttention}(Q,[\mathbf{F}_{i-1},\mathbf{F}_{i},\mathbf{F}_{i+1}])$ (2)
-Similar to previous methods [26, 8], a projection layer is needed to align the features of vision encoder and language models.
-However, we observe that utilizing a single projection layer for both the visual and text-related tokens may lead to subpar performance (more discussions ).
-Since the two types of features from the visual compressor and the textual region compressor are regarded as different representations of one frame, we employ two projection layers to transform $\mathbf{F}^{v}_{i}$ and $\mathbf{F}^{t}_{i}$ into $\hat{\mathbf{F}}^{v}_{i}$ and $\hat{\mathbf{F}}^{t}_{i}$, respectively.
-Finally, the global visual tokens ^$\hat{\mathcal{F}}^{v}$ and the text-related tokens ^$\hat{\mathcal{F}}^{t}$ of all frames are interleaved into the output of InterleavedVT $\mathbf{F}^{out}$:
-$\mathbf{F}^{out}=[\hat{\mathbf{F}}^{v}_{0},\hat{\mathbf{F}}^{t}_{0},...,\hat{\mathbf{F}}^{v}_{N-1},\hat{\mathbf{F}}^{t}_{N-1}]$ (3)
-Input : Predicted subtitles $\mathcal{S}$={$S_{0}$, $S_{1}$, …, $S_{T-1}$}, where ; OCR results of all frame in the input video $\mathcal{O}$={$O_{0}$, $O_{1}$, …., $O_{L-1}$}, where $O_{l}$ = $\text{o}_{0}$, $\text{o}_{1}$,…, $\text{o}_{K_{l}-1}$}; Similarity threshold $\mathbf{sim}$, Raw FPS of the input video $\mathbf{f}$, Subtile matching range $\mathbf{r}$
-Output : Subtitles in SRT format $\mathcal{S}^{\prime}$={$S^{\prime}_{0}$, $S^{\prime}_{1}$, …, $S^{\prime}_{T-1}$}, $S^{\prime}_{t}$=[$b^{\prime}_{t}$, $e^{\prime}_{t}$, $s^{\prime}_{t}$]
+III-C Spatiotemporal Subtitle-Salient Module
+The proposed S3 module comprises the following two branches:
+III-C1 Spatial Semantic Context Aggregate
+Subtitle generation is not a pure text-recognition problem but is inherently guided by the visual semantics of each frame.
+Movie and short-video subtitles often correspond to high-level scene cues—such as character interactions, camera composition, and background layout—that provide strong prior information about whether subtitles appear and how they should be phrased.
+To capture such global semantics in an efficient and robust manner, we introduce the Spatial Semantic Context Aggregate (SSCA) branch.
+SSCA summarizes the frame feature $\mathbf{F}_{i}$ into a compact layout-aware representation $\mathbf{F}^{v}_{i}\in\mathbb{R}^{p\times p\times C}$ using spatial average pooling:
+$\mathbf{F}^{v}_{i}=\texttt{AvgPool}(\mathbf{F}_{i}).$ (1)
+Although the operation is simple, it serves a clear and task-aligned role: it suppresses low-level background noise while preserving high-level semantic structure—such as scene geometry, speaker regions, and global motion cues—that influence subtitle presence and content.
+SSCA thus provides a stable global prior that complements the text-focused modeling of the TSTQ branch.
+III-C2 Temporal Subtitle Token Query
+While SSCA captures holistic scene cues, accurate subtitle extraction further requires identifying fine-grained visual evidence related to on-screen text.
+To this end, we adopt a lightweight query-based design inspired by Q-Former [li2023blip].
+A set of learnable query embeddings $\mathbf{Q}$ are used as latent probes that attend to subtitle-salient regions in the dense feature map.
+Each query specializes in capturing complementary cues—such as text contours, stroke patterns, or high-contrast regions—by performing cross-attention with $\mathbf{F}_{i}$:
+$\mathbf{F}^{t}_{i}=\texttt{CrossAttention}(\mathbf{Q},\mathbf{F}_{i}).$ (2)
+Unlike general vision–language alignment, subtitle extraction requires strong temporal stability because subtitles usually persist across several frames and gradually fade in or out.
+To enhance robustness, we extend the attention context to a temporal window consisting of the current frame and its neighbors:
+$\mathbf{F}^{t}_{i}=\texttt{CrossAttention}(\mathbf{Q},[\mathbf{F}_{i-1},\mathbf{F}_{i},\mathbf{F}_{i+1}]).$ (3)
+This temporal aggregation enables TSTQ to track slowly transitioning subtitle patterns, reduce frame-wise inconsistencies, and generate subtitle-related tokens that are both temporally consistent and visually discriminative.
+III-C3 Feature Alignment and Fusion
+The SSCA and TSTQ branches provide two complementary perspectives on subtitle perception: SSCA encodes global spatial semantics, while TSTQ extracts localized, temporally coherent subtitle tokens.
+Although both originate from the same visual encoder, their feature distributions vary substantially—one emphasizes holistic context, and the other emphasizes fine-grained text evidence.
+To align these heterogeneous representations with LLM, we introduce the separate projector (SP) strategy :
+$\hat{\mathbf{F}}^{v}_{i}=\texttt{Proj}_{v}(\mathbf{F}^{v}_{i}),\quad\hat{\mathbf{F}}^{t}_{i}=\texttt{Proj}_{t}(\mathbf{F}^{t}_{i}).$ (4)
+The separate projector strategy prevents feature interference and preserves the modality-specific characteristics essential for subtitle reasoning.
+Finally, the projected global semantics and subtitle-salient tokens across all frames are interleaved to form the unified output sequence:
+$\mathbf{F}^{out}=[\hat{\mathbf{F}}^{v}_{0},\hat{\mathbf{F}}^{t}_{0},...,\hat{\mathbf{F}}^{v}_{N-1},\hat{\mathbf{F}}^{t}_{N-1}].$ (5)
+This fused representation integrates both spatial and temporal cues in a subtitle-aware manner, enabling the language model to produce temporally aligned and semantically faithful subtitles.
+Input : Predicted subtitles $\mathcal{S}$$S_{0}$, $S_{1}$, …, $S_{T-1}$}; OCR results of all frame in the input video $\mathcal{O}$$O_{0}$, $O_{1}$, …., $O_{L-1}$}, where $O_{l}$ = $\text{o}_{0}$, $\text{o}_{1}$,…, $\text{o}_{K_{l}-1}$}; Similarity threshold $\mathbf{sim}$; Raw FPS of the input video $\mathbf{f}$; Subtile matching range $\mathbf{r}$.
+Output : Subtitles in SRT format $\mathcal{S}^{\prime}$$S^{\prime}_{0}$, $S^{\prime}_{1}$, …, $S^{\prime}_{T-1}$}, $S^{\prime}_{t}$$b^{\prime}_{t}$, $e^{\prime}_{t}$, $s^{\prime}_{t}$]
+1ex
 foreach subtitle $S$ in $\mathcal{S}$ do
@@
   # perform matching in the range $\mathbf{r}$ to locate the accurate start timestamp ;
-       for $i$=$bc-\mathbf{r}$ to $b_{c}+\mathbf{r}$ do
+[…]
      $min_{\text{ed}}$ = $\infty$;
@@
 […]
-                         $min_{\text{ed}}$ = EditDistance(s, o);
+           $min_{\text{ed}}$ = EditDistance(S,o)(S,o);
          end if
@@
 return Subtitles in SRT format $\mathcal{S}^{\prime}$ ;
-Algorithm 1 Convert EVE predictions into subtitles in SRT format.
-Large Language Model.
-In the past few years, numerous foundation large language models are open-sourced, such as Qwen [40] and LLaMa [29].
-Since the proposed EVE aims at solving bilingual (i.e., English and Chinese) video subtitle extraction, we select Qwen as the large language model in EVE.
+Algorithm 1 Converting EVE predictions into subtitles with SRT format.
+III-D Large Language Model
+In the past few years, numerous foundation large language models are open-sourced, such as Qwen2 [yang2024qwen2] and LLaMa [touvron2023llama].
+Since the proposed EVE aims at solving bilingual (i.e., English and Chinese) video subtitle extraction, we select Qwen2 as the large language model of EVE.
 In addition, taking into account the time efficiency in practical applications, we build EVE-7B and EVE-0.5B based on Qwen2-7B and Qwen2-0.5B, respectively.
-Detail performance comparison is shown in Sec.
-5.
-3.2 Training
-The training process of the proposed EVE is divided into two stages: pre-training stage and supervised fine-tuning stage. We compare these two different stages from the following three aspects:
+Detail performance comparison is shown in Section V.
+III-E Training
+The training process of the proposed EVE consists of two stages: a pre-training stage and a supervised fine-tuning stage. We describe these two stages in detail from the following three aspects:
   • Training Details.
-The pre-training stage primarily aims at enhancing the text perception capability of EVE, enabling it to predict all text tracklets in video frames.
-As shown in Figure 3, at the pre-training stage, the large language model is frozen while the vision encoder and the proposed adapter InterleavedVT are optimized.
-After the pre-training stage, the vision encoder is well trained to perceive texts.
-Thus, at the supervised fine-tuning stage, the vision encoder is frozen while the adapter InterleavedVT and the language model are optimized to learn discriminate whether texts in the input video are subtitles.
+The pre-training stage aims to enhance the text perception capability of EVE by enabling it to predict all text tracklets appearing in video frames.
+As shown in Figure 3, during this stage, the large language model is frozen, while the vision encoder and the proposed S3 module are jointly optimized.
+After pre-training, the vision encoder becomes proficient at perceiving textual regions.
+Consequently, in the supervised fine-tuning stage, the vision encoder is frozen, and only the S3 module together with the large language model are trained to distinguish whether the detected texts correspond to subtitles.
   • Input.
-Both pre-training and supervised fine-tuning stages share the same input template.
-In the template, we additionally prepend the corresponding frame index to the visual tokens to improve the temporal perception ability of our model.
-The limit for the number of input frames is set to 30 during pre-training while it is set to 120 during supervised fine-tuning.
+Both stages share the same input template.
+To improve temporal perception, we prepend the corresponding frame index to each frame’s visual tokens.
+The number of input frames is limited to 30 during pre-training and 120 during supervised fine-tuning.
   • Prediction.
-For prediction, the two stages also share the same template, where the start/end frame index and texts are included.
-Differently, the pre-training stage learns to recognize all text tracklets in videos while the supervised fine-tuning stage only considers those subtitle tracklets.
-Figure 3: Training details of EVE. The LLM is frozen during pre-training while the vision encoder is forzen during supervised fine-tuning.
-3.3 Inference
+The prediction template is also identical across the two stages, containing the start frame index, end frame index, and texts.
+The difference lies in their objectives: the pre-training stage learns to recognize all text tracklets in videos, whereas the supervised fine-tuning stage focuses solely on those corresponding to subtitles.
+Figure 3: Training details of EVE. The LLM is frozen during pre-training while the vision encoder is frozen during supervised fine-tuning.
+III-F Inference
+III-F1 Output Format
+We use the SRT files as the model’s output, as they are one of the most popular subtitle file formats for video content.
+An SRT file stores sequential subtitle texts with corresponding timestamps but no audio or video data, resulting in a compact file size.
+In addition, the SRT file needs to specify the line number, subtitle content, when the text should appear on the video, and when it should disappear to make room for the next line, as shown in Figure 4.
+III-F2 Frame Sampling and Post-processing
 Due to the limitation of EVE in input token length, inference cannot be completed at high frame rates.
@@
 Although the predicted frame index can directly correspond to a precise timestamp (i.e., hour, minute, and second), it does not represent the precise time at which each subtitle appears and disappears in the video.
-To output more accurate subtitle content, such as in SRT format (examples are provided in the supplementary materials), we propose a post-processing method based on OCR results at full frame rate.
+To output more accurate subtitle content, such as in SRT format, we propose a post-processing method based on OCR results at full frame rate.
 The algorithm of post-processing is described in Algorithm 1.
@@
 Finally, by calculating the edit distance between predicted texts and those within adjacent frames, the precise subtitle timestamp can be obtained.
+Example of a bilingual SRT subtitle file 1 00:00:01,739 $\rightarrow$ 00:00:02,826 那么我是谁?
+Then who am I?
+2 00:00:03,521 $\rightarrow$ 00:00:04,913 你就是那个节目的明星 You’re the star.
+5 00:00:04,540 $\rightarrow$ 00:00:05,826 所以才有那么多人看你 That’s what made you so good to watch.
+Figure 4: An example of a bilingual SRT subtitle file.
 It is worth mentioning that the aforementioned post-processing is not necessary for EVE.
@@
 Therefore, it can be said that EVE is an end-to-end method for video subtitle extraction.
+TABLE I: The number of samples in different languages. Both short videos and movie videos are mainly in Chinese.
 Chinese English Bilingual None-Text
@@
 Movie 3,141 2,600 173 5,258
-Table 1: The number of samples in different languages. Both short videos and movie videos are mainly in Chinese.
-4 Data Construction
-In this paper, we propose a large dataset, called ViSa, to benchmark the video subtitle extraction task.
-The raw videos are collect from two datasets: CNVid [9] and MovieChat-1K [27].
+IV Dataset
+Although video subtitle extraction has been studied for years, existing datasets for this task remain extremely limited.
+To this end, we present the ViSa dataset, the first large-scale dataset for video subtitle extraction, providing training material and a evaluation benchmark dataset for this task.
+In accordance with the previously mentioned training objectives, we divide ViSa into three components: the pre-training dataset, the supervised fine-tuning dataset, and the test dataset.
+IV-A Comparison With Existing Dataset
+Elshahaby et al.
+[elshahaby2022end] introduced the FiViD dataset for video subtitle extraction task, while Yan et al.
+[yan2020end] trained and evaluated their model on publicly available image-based OCR datasets, including MJ Synth [jaderberg2014synthetic], ICDAR03 [1227749], ICDAR13 [6628859], and SVT [Wang20111457].
+In contrast, Mathur et al.
+[mathur2015generating] relied on the audio-only VoxForge [voxforge] corpus to explore speech-to-text subtitle generation, a direction that falls outside the scope of the visual subtitle extraction task studied in this work.
+As shown in Table  II, the aforementioned datasets are typically limited in scale and lack timestamp annotations.
+Moreover, most of them composed of images rather than video.
+Consequently, these datasets are insufficient to support modern LVLM-based models that require large-scale, high-quality supervision, especially for fine-grained temporal localization.
+Meanwhile, recent large video datasets such as CNVid [gan2023cnvid] and MovieChat-1K [song2024moviechat] provide massive video–text pairs, yet they are not designed for subtitle extraction.
+They contain video captions or movie subtitles that (1) do not align with frame-level timestamps, (2) are not guaranteed to reflect on-screen subtitle content, and (3) lack start–end boundaries required by this task.
+Thus, they cannot be directly used to train or evaluate a subtitle extraction model.
+To fill this gap, we construct ViSa, the first large-scale dataset specifically designed for timestamped video subtitle extraction.
+ViSa contains bilingual (Chinese and English) subtitle texts paired with both coarse-grained and fine-grained timestamps, enabling scalable pre-training, supervised fine-tuning, and rigorous evaluation of video subtitle extraction systems.
+Table II summarizes the differences between ViSa and prior subtitle extraction datasets.
+TABLE II: Comparison of ViSa with datasets used in previous works.
+Dataset
+Data
+Type
+Scale
+On-screen Text Annot.
+Time-
+stamp
+FiViD [elshahaby2022end] Image $\approx$74k Yes No
+MJ Synth [jaderberg2014synthetic] Synthetic 9M Yes No
+ICDAR03 [1227749] Image 860 Yes No
+ICDAR13 [6628859] Image 857 Yes No
+SVT [Wang20111457] Image 647 Yes No
+CNVid [gan2023cnvid] Video 3.5M No No
+MovieChat-1K [song2024moviechat] Video 1k No No
+ViSa (ours) Video 2.5M Yes Yes
+IV-B Data Construction
+The raw videos of ViSa are collected from two datasets: CNVid [gan2023cnvid] and MovieChat-1K [song2024moviechat].
 CNVid is the largest public video-text dataset in Chinese, which contains 3.5M short videos collect from Douyin  [^6] https://www.douyin.com/.
@@
 For CNVid, we perform the following three steps to reserve the ideal videos: (1) First, we filter out videos with a duration shorter than 10 seconds or longer than 120 seconds.
-The reason for this operation is that short videos (duration$<$10s) contain a lot of noise while the number of frames in long videos (duration$>$120s) exceeds the model’s input limits.
+The reason for this operation is that short videos (duration<<10s) contain a lot of noise while the number of frames in long videos (duration>>120s) exceeds the model’s input limits.
 (2) Then, we utilize a in-house video OCR tools to predict the text tracklets in the reserved videos.
@@
 The proportion of videos without subtitles in movie videos is higher that in short videos.
-Since we clip movie videos based on a uniformly sampled duration in the range of 15--60 seconds, the duration distribution of movie videos is more balanced compared to short videos.
-Specifically, ViSa divides the collected videos into three parts: pre-training, supervised fine-tuning, and test datasets.
-Some statistics of them are described in the following.
-More details are shown in the supplementary materials.
-Pre-training dataset.
+Since we clip movie videos based on a uniformly sampled duration in the range of 15-60 seconds, the duration distribution of movie videos is more balanced compared to short videos.
+Specifically, ViSa divides the collected videos into three parts: pre-training, supervised fine-tuning, and test datasets. Some statistics of them are described below.
+IV-C Pre-training dataset.
 The pre-training dataset contains 2.35M videos, all of which are collected from CNVid.
 The videos in the pre-training dataset are annotated with text tracking and text recognition since the goal of pre-training is to equip the model with the text perception capability.
-The statistics of pre-training dataset are shown in the supplementary materials.
+Therefore, this annotation is coarse-grained and low-cost, considering the large scale of this portion of the data.
 The majority of pre-training videos are typically shorter than 60 seconds in length.
 The number of text tracklets in videos mostly ranges between 12 and 30.
-In a total of 2.5M pre-training videos, 1.8M videos include text tracklets while 0.7M videos contain no texts.
-Supervised Fine-tuning (SFT) dataset.
-The videos for supervised fine-tuning originate from both CNVid and MovieChat-1K.
-Concretely, the SFT dataset includes 210K Chinese videos from CNVid and 20K bilingual videos from MovideChat-1K.
+In a total of 2.35M pre-training videos, 1.8M videos include text tracklets while 0.7M videos contain no texts.
+IV-D Supervised Fine-Tuning dataset.
+The videos for Supervised Fine-Tuning (SFT) originate from both CNVid and MovieChat-1K.
+Concretely, the SFT dataset includes 210K Chinese videos from CNVid and 20K bilingual videos from MovieChat-1K.
 Each video in the SFT dataset is annotated with subtitle text tracking and corresponding results.
+Specifically, we use the annotation tool arctime  [^7] https://arctime.org to perform automated annotation for each video, followed by manual review and correction to ensure the subtitles are accurate and the timestamp error is less than 0.5 seconds.
 As a data augmentation strategy to expand the SFT dataset, we annotate each video with both 1-FPS and 2-FPS sampling rate.
-The statistics of SFT dataset are shown in Fig.
-4.
-As shown in Tab.
-1, among 230K SFT videos, 30K videos contains no texts.
-Test dataset.
+The statistics of SFT dataset are shown in Figure 5.
+As shown in Table I, among 230K SFT videos, 30K videos does not contain text.
+IV-E Test dataset.
 To evaluate the performance of EVE on both short videos and movies, we collect 857 videos from CNVid and 862 videos from MovieChat-1K to build the test dataset.
-The annotations of this dataset are the same as that of the SFT dataset and they are manually annotated.
+The annotation for this portion of the data is fine-grained and manually generated to enable objective model evaluation.
 Similar to the SFT dataset, both positive samples (i.e., videos containing texts) and negative samples (i.e., videos containing no texts) are also included in the test dataset.
+TABLE III: Performance comparison with existing three types of methods.
 Method Short Video Movie
@@
 NED SubER SubER NED SubER SubER
-Open-sourced Tools VideOCR [4] 0.348 72.72 52.63 0.686 6.28 57.12
-VideOCR [4] + Subregion 0.659 52.92 43.81 0.702 5.24 66.35
-PyVideoTrans [3] 0.477 46.03 92.34 0.639 88.35 52.51
-Cloud Service OpenVision [2] 0.771 21.85 42.85 0.718 6.57 18.61
-GhostCut [1] 0.655 54.78 46.00 0.659 7.26 24.43
-Open-sourced LVLMs Qwen2-VL-2B [31] 0.068 - - 0.070 - -
-Qwen2-VL-7B [31] 0.123 - - 0.227 - -
-MiniCPM-V-2.6 [42] 0.043 - - 0.047 - -
+Open-sourced Tools VideOCR [videocr] 0.348 72.72 52.63 0.686 6.28 57.12
+VideOCR [videocr] + Subregion 0.659 52.92 43.81 0.702 5.24 66.35
+PyVideoTrans [pyvideotrans] 0.477 46.03 92.34 0.639 88.35 52.51
+Cloud Service OpenVision [openvision] 0.771 21.85 42.85 0.718 6.57 18.61
+GhostCut [ghost] 0.655 54.78 46.00 0.659 7.26 24.43
+Open-sourced LVLMs Qwen2-VL-2B [wang2024qwen2] 0.068 - - 0.070 - -
+Qwen2-VL-7B [wang2024qwen2] 0.123 - - 0.227 - -
+MiniCPM-V-2.6 [yao2024minicpm] 0.043 - - 0.047 - -
 Ours EVE-0.5B 0.759 22.82 18.36 0.685 5.10 17.06
 EVE-7B 0.882 15.16 13.92 0.808 1.64 10.15
-Table 2: Performance comparison with existing three types of methods.
-Since the open-sourced LVLMs are trained with no video subtitle extraction datasets, they cannot achieve ideal performance on this task.
-5 Experiments
-5.1 Implementation Details
+V Experiments
+V-A Implementation Details
 The proposed method is implemented with PyTorch.
 All experiments are conducted on $24$ A100 GPUs.
-The proposed EVE is trained using the AdamW optimizer [20] where the learning rate is set to $5\times 10^{-5}$.
+The proposed EVE is trained using the AdamW optimizer [loshchilov2017decoupled] where the learning rate is set to $5\times 10^{-5}$.
 The input frames are resized into $640\times 640$.
 The batch size is set to 48 for pre-training and 16 for supervised fine-tuning.
-The number of queries $K$ in the textual region compressor is set to 10, and the number of visual tokens is set to 16.
-More relevant experimental results are shown in the supplementary materials.
-The vision encoder is initialized with the pre-trained parameters of Donut [11], and the large language model loads the pre-trained models of Qwen2-7B or Qwen2-0.5B.
-5.2 Metrics
-In this paper, we adopt two types of metrics to evaluate the performance of models: Normalized Edit Distance (NED) [44] and SubER [33]. NED is defined as:
+The shape of $\mathbf{F}^{v}_{i}$ is set to $\mathbb{R}^{4\times 4\times C}$.
+The number of queries $Q$ in the TSTQ is set to 10, and the number of visual tokens after SSCA is set to 16.
+The vision encoder is initialized with the pre-trained parameters of Donut [kim2021donut], and the large language model loads the pre-trained models of Qwen2-7B or Qwen2-0.5B.
+V-B Metrics
+In this paper, we adopt two types of metrics to evaluate the performance of models: Normalized Edit Distance (NED) [zhang2019icdar] and SubER [wilken2022suber]. NED is defined as:
 $\texttt{NED}(\hat{\mathbf{y}}_{i},\mathbf{y}_{i})=1-\frac{1}{S}\sum^{S}_{i=1}\texttt{ED}(\hat{\mathbf{y}}_{i},\mathbf{y}_{i})/\texttt{MaxL}(\hat{\mathbf{y}}_{i},\mathbf{y}_{i})$ (6)
@@
 However, it may hardly evaluate the accuracy of timestamp predictions.
-Therefore, we additionally adopt SubER [33] to evaluate the performance of models.
+Therefore, we additionally adopt SubER [wilken2022suber] to evaluate the performance of models.
 SubER is defined as:
@@
 word edits are insertions, deletions and substitutions of words; break edits are insertions, deletions and substitutions of line breaks; shifts are movements of one or more adjacent hypothesis tokens to a position of a matching phrase in the reference; reference words/reference breaks is the ground-truth words/breaks.
-More details about SubER can refer to [33].
-5.3 Experimental Results
-Evaluating Existing Methods.
-Although the video subtitle extraction task has been explored before, there is still a lack of comprehensive dataset to benchmark this task.
-In this paper, we propose a video subtitle extraction dataset containing pre-training, supervised fine-tuning and test datasets.
-We have evaluated three categories of video subtitle extraction methods on the proposed test dataset.
-Among the open-sourced tools, PyVideoTrans based on ASR achieves relatively inferior performance on negative samples compared with VideOCR.
-The reason for this results is that the audio in the negative sample will also be translated into sentences, which are regarded as subtitles of the input video.
-We have also evaluated two cloud services that can perform video subtitle extraction.
-Ghost cannot perform well on short videos containing a lot of background texts since it tends to mistake background texts as subtitles.
+More details about SubER can refer to [wilken2022suber].
+TABLE IV: Performance comparison with two typical token compression methods.
+Method Short Video Movie
+Positive Negative Positive Negative
+NED SubER SubER NED SubER SubER
+T-Selector [wang2025elysium] 0.801 20.32 23.84 0.731 8.72 19.30
+ToMe [bolyatoken] 0.714 28.35 25.62 0.705 10.79 15.59
+EVE-0.5B 0.759 22.82 18.36 0.685 5.10 17.06
+EVE-7B 0.882 15.16 13.92 0.808 1.64 10.15
+V-C Experimental Results
+V-C1 Evaluating Existing Methods
+We evaluated three categories of video subtitle extraction methods on the proposed test dataset as shown in Table  III.
+Among the open-sourced tools, PyVideoTrans based on Automatic Speech Recognition (ASR) achieves relatively inferior performance on negative samples compared with VideOCR.
+The reason for these results is that ASR-based methods inherently treat spoken audio as potential subtitles, causing false positives in negative samples where no on-screen text exists.
+We also evaluated two cloud services that can perform video subtitle extraction.
+GhostCut cannot perform well on short videos containing a lot of background texts since it tends to mistake background texts as subtitles.
 In contrast, OpenVision can achieve relatively ideal performance on both short videos and movie videos.
-Two open-source multi-modal large models (Qwen2-VL [31] and MiniCPM-V-2.6 [42]) perform poorly on the video subtitle extraction task because they are not trained on corresponding datasets.
+Two open-source multi-modal large models (Qwen2-VL [wang2024qwen2] and MiniCPM-V-2.6 [yao2024minicpm]) perform poorly on the video subtitle extraction task because they are not trained on corresponding datasets.
 They prefer to performing video captioning rather than extracting subtitle texts, thus resulting in subpar performance.
 In addition, the experimental results demonstrate that Qwen2-VL’s video OCR capability is better than MiniCPM-V-2.6.
-More visualizations are shown in supplementary materials.
-Comparison with Existing Methods.
-In this paper, we establish two models with different LLMs.
-To achieve better performance in video subtitle extraction, we utilize Qwen2-7B to construct EVE-7B.
-Through the experimental results in Tab.
-2, EVE-7B achieves the best performance compared with three types of video subtitle extraction methods.
-For example, on the short video dataset, EVE-7B outperforms OpenVision by 11.1% in the metric of NED.
-We have visualized an example in Fig 5, and the result shows that the proposed EVE can well predict the subtitles without missing and repetition problem.
-More visualizations are shown in the supplementary materials.
-We have also conducted experiments to compare with two typical token compression methods (T-Selector [30] and ToMe [5]).
-The results are shown in the supplementary materials.
-Considering that the EVE-7B will cost more computing resources in practical applications, we establish a light-weighted model EVE-0.5B with Qwen2-0.5B.
-The experimental results in Tab.
-2 demonstrate that despite using the same dataset for training, EVE-0.5B still cannot achieve the same performance as EVE-7B.
-However, benefiting from the proposed InterleavedVT and corresponding dataset ViSa, it can still surpass most of existing methods.
-Figure 5: An example of subtitle extraction for different methods.
-The ASR based method PyVideoTrans tends to miss some audio-irrelative texts; Qwen2-VL performs video captioning rather than extracting subtitles; OpenVision may be based on text tracking, resulting repetition problem when text tracklets are wrongly predicted.
-Gray and Red indicate the missing and repetition prediction problem, respectively.
-Green represents the format of the answer is incorrect.
-VC TRC SP Short Video Movie
+TABLE V: Ablation studies on the proposed EVE (metric: SubER).
+SSCA TSTQ SP Short Video Movie
 Pos Neg Pos Neg
@@
 15.16 13.92 1.64 10.15
-Table 3: Ablation studies on the proposed EVE (metric: SubER). Through experimental results, the separate projector is necessary for InterleavedVT.
-5.4 Ablation Studies
-In this paper, we propose an innovative adapter called InterleavedVT, which includes a visual compressor (VC) and a textual region compressor (TRC) for compressing visual features.
-Additionally, we introduce the separate projector (SP) to project the visual tokens and text-related tokens into the input space of large language models.
+V-C2 Comparison with Existing Methods
+In this paper, we establish two models with different LLMs.
+To achieve better performance in video subtitle extraction, we utilize Qwen2-7B to construct EVE-7B.
+Through the experimental results in Table III, EVE-7B achieves the best performance compared with three types of video subtitle extraction methods.
+For example, on the short video dataset, EVE-7B outperforms OpenVision by 11.1% in the metric of NED.
+Considering that the EVE-7B will cost more computing resources in practical applications, we establish a light-weighted model EVE-0.5B with Qwen2-0.5B.
+The experimental results in Table III demonstrate that despite using the same dataset for training, EVE-0.5B still cannot achieve the same performance as EVE-7B.
+However, benefiting from the proposed framework and training strategy, it can still surpass most of existing methods.
+Figure 6: An example of subtitle extraction for different methods.
+Gray and Yellow indicate the missing and repetition prediction problem, respectively.
+Red represents the format of the answer is incorrect.
+Green represent the correct prediction content.
+Figure 7: The visualization for the temporal subtitle token querying using heatmap.
+Figure 8: Visualization of subtitle extraction results with timestamps from a movie clip using MiniCPM, GhostCut, VideOCR, Qwen2-VL, and the proposed EVE.
+V-D Comparison with Token Compression Methods
+Beyond comparisons with existing subtitle extraction tools and LVLM-based baselines, we further evaluate the proposed S3 module against two representative token compression approaches, T-Selector [wang2025elysium] and ToMe [bolyatoken], as summarized in Table IV.
+As mentioned in the Related Work section, these methods are functionally similar to the proposed S3 module, as they all aim to extract key information to feed into the LLM.
+Therefore, we conduct this experiment for comparison.
+Both methods are re-implemented using the Qwen2-7B backbone to ensure a fair comparison under identical conditions.
+The results reveal that despite utilizing a smaller 0.5B language model, EVE-0.5B achieves performance on par with or even surpasses these token compression baselines, validating the efficiency of our design.
+When scaled to EVE-7B, the performance improves substantially across all benchmarks.
+For instance, compared to ToMe, EVE achieves a 16.8% improvement in NED on the Short Videos test set, demonstrating stronger temporal alignment and finer-grained token utilization.
+This consistent advantage highlights the effectiveness of our proposed S3 module, which interleaves vision and text representations to selectively preserve semantically relevant tokens while filtering out redundant ones.
+Unlike conventional compression schemes that rely purely on token similarity, S3 module adaptively integrates multimodal cues, enabling more precise identification of subtitle-relevant regions and enhancing overall extraction robustness across both short and long-form videos.
+V-E Ablation Studies
+In this paper, we propose an innovative adapter called S3 module, which includes a Spatial Semantic Context Aggregate (SSCA) and a Temporal Subtitle Token Query (TSTQ).
+Additionally, we introduce the separate projector (SP) to project the global visual tokens and text-related tokens into the input space of large language models.
 In this section, we conduct ablation studies on these three modules.
-The experimental results, as shown in Tab.
-3, indicate that introducing either the visual compressor or the textual region compressor alone does not outperform previous methods.
-However, when both of them are integrated into InterleavedVT, there is a significant improvement in model performance.
-Compared to only using the textual region compressor, our method achieves a 5.03% and 5.28% improvement in SubER metrics on the positive samples of short video and movie datasets, respectively.
+The experimental results, as shown in Table V, indicate that introducing either the SSCA or the TSTQ alone does not outperform previous methods.
+However, when both of them are integrated into S3 module, there is a significant improvement in model performance.
+Compared to only using the TSTQ, our method achieves a 5.03% and 5.28% improvement in SubER metrics on the positive samples of short video and movie datasets, respectively.
 Furthermore, introducing separate projectors can further improve the performance.
 A plausible explanation is that the two types of information focus on different aspects of visual effects, and thus, separate projectors facilitate better alignment of features with LLMs.
-6 Limitations and Future Work
-Although the proposed EVE have compressed each frame into 26 tokens and supported to input 120 frames for video subtitle extraction, it still have difficulties in solving long videos (e.g., longer than one hour).
-In future work, we will explore to more effectively compress text-related tokens in video frames to increase the number of image inputs to the model.
-Furthermore, the proposed EVE can only predict timestamps and texts of subtitles, ignoring the positional information of texts (i.e., bounding box).
-To support more downstream applications, we will attempt to predict the accurate coordinates of subtitles in each frame.
-7 Conclusion
-In this paper, we propose an End-to-end Video Subtitle Extraction model, called EVE, which consists of a vision encoder, an adapater module and a large language model.
-In EVE, a novel adapter, called InterleavedVT, is introduced for effectively compressing visual tokens.
+V-F Hyperparameter Selection
+There are two hyperparameters in EVE: the number of queries in the TSTQ and the number of visual tokens after SSCA.
+We have conducted corresponding experiments with different numbers of tokens.
+When the number of visual tokens (denoted as $N$) and the number of queries in the TSTQ (denoted as $K$) are set to larger values of (NN=36 and KK=20), the performance of EVE is only improved by 0.3% in the NED metric, but the computational cost will increase significantly, due to more tokens being reserved for the LLM.
+When $N$ and $K$ are set to smaller values (NN=4 and KK=1), there is a significant decrease in the performance of EVE, with the NED metric decreasing by nearly 28.6%.
+In order to balance computational cost and model performance, we ultimately set $N$ and $K$ to 16 and 10, respectively.
+V-G Visualization for Temporal Subtitle Token Query
+To better understand the role of the Temporal Subtitle Token Query in EVE, we visualize its attention responses across representative frames, as shown in Figure 7.
+The upper row depicts the original video frames, while the lower row illustrates the corresponding attention heatmaps.
+We observe that the TSTQ consistently activates in regions containing textual elements, such as interface buttons, titles, and overlaid text cues.
+This selective activation indicates that the module effectively distinguishes text-related areas from irrelevant visual content, enabling more efficient information compression and representation.
+Moreover, as the video progresses, the attention maps dynamically adjust to newly appearing text, demonstrating strong adaptability to temporal variations.
+These visualizations validate that the TSTQ successfully enhances the model’s perception of text-bearing regions and contributes to more precise multimodal understanding within EVE.
+V-H Visualization for Video Subtitle Extraction
+In Figure 6, we compare PyVideoTrans [pyvideotrans], OpenVision [openvision], Qwen2-VL [wang2024qwen2], and the proposed EVE for subtitle extraction.
+PyVideoTrans omits critical dialogue segments, while OpenVision introduces redundancy through repeated outputs.
+Qwen2-VL diverges from the task objective, producing scene-level descriptions rather than precise transcripts.
+By contrast, EVE generates temporally aligned, semantically faithful, and duplication-free subtitles.
+For example, the utterance “Oh, have they gone off again, have they?” is recovered accurately and uniquely.
+These results confirm EVE’s robustness and highlight its effectiveness as a reliable framework for multimodal subtitle generation.
+In Figure 8, we further compare MiniCPM  [yao2024minicpm], GhostCut [ghost], VideOCR [videocr], Qwen2-VL [wang2024qwen2], and EVE for video subtitle extraction.
+MiniCPM fails to perform transcription and instead outputs verbose scene-level descriptions.
+GhostCut captures most subtitle lines but exhibits minor temporal misalignment and redundant repetitions.
+VideOCR shows high recognition quality but suffers from token concatenation issues such as “It’s ok, Truman.
+I understand.” caused by inaccurate text segmentation.
+Qwen2-VL generates only a single incomplete line, failing to cover the entire dialogue sequence.
+In contrast, EVE delivers temporally consistent, structurally complete, and duplication-free subtitles.
+For instance, the utterance “It’s ok, Truman.
+I understand.” is precisely aligned and clearly formatted, demonstrating EVE’s strong capability in maintaining accurate temporal correspondence and structural completeness.
+Moreover, the visualization results indicate that the timestamps predicted by EVE are closer to the ground truth compared to other methods.
+VI Conclusion
+In this paper, we propose an End-to-end Video Subtitle Extraction model, called EVE, which consists of a vision encoder, an adapter module and a large language model.
+In EVE, a novel adapter called the S3 module has been introduced, which is able to model salient subtitle representations while retaining only a minimal number of tokens.
 To benchmark the video subtitle extraction task, we propose a large bilingual (i.e., Chinese and English) dataset ViSa, including 2.35M videos for pre-training, 230K videos for fine-tuning and 1.6K videos for test.
 Through experimental results, we observe that the proposed EVE achieves the best performance in both NED and SubER compared with previous methods for video subtitle extraction.
-Supplementary Material
-8 SRT Format
-An SRT file is one of the most popular subtitle file formats for video content.
-These plain text files include the text of the subtitles in sequence, along with the start and end timestamps.
-However, an SRT file does not contain any video or audio content.
-That’s why the file sizes are so small.
-An SRT file needs to specify the line number, subtitle content, when the text should appear on the video, and when it should disappear to make room for the next line.
-An example is shown in the following.
-Figure 6: An example of SRT files.
-9 Hyperparameter Selection
-There are two hyperparameters in EVE: the number of queries in the textual region compressor and the number of visual tokens.
-We have conducted corresponding experiments with different numbers of tokens.
-When the number of visual tokens (denoted as $N$) and the number of queries in the textual region compressor (denoted as $K$) are set to larger values of ($N$=36 and $K$=20), the performance of EVE is only improved by 0.3% in the NED metric, but the computational cost will increase fourfold.
-When $N$ and $K$ are set to smaller values ($N$=4 and $K$=1), there is a significant decrease in the performance of EVE, with the NED metric decreasing by nearly 28.6%.
-In order to balance computational cost and model performance, we ultimately set $N$ and $K$ to 16 and 10, respectively.
-10 Comparison with Token Merging Methods
-Except the open-sourced subtitle extraction tools and LVLMs, we also compare our method with two typical token Merging methods (T-Selector [30] and ToMe [5]).
-The results are shown in Tab.
-4.
-We use Qwen2-7B to implement T-Selector and ToMe for subtitle extraction.
-The experimental results show that although our method EVE-0.5B only adopts a 0.5B language model, it can achieve comparable results to T-Selector and ToMe.
-When we use Qwen2-7B for EVE, our method can achieve significant performance improvements.
-For example, compared to ToMe , EVE can achieve a 16.8% performance improvement (in NED) on the Short Videos test dataset.
-The results demonstrate that the proposed token merging method InterleavedVT can more efficiently extract the subtitle-related information in videos.
-Method Short Video Movie
-Positive Negative Positive Negative
-NED SubER SubER NED SubER SubER
-T-Selector [30] 0.801 20.32 23.84 0.731 8.72 19.30
-ToMe [5] 0.714 28.35 25.62 0.705 10.79 15.59
-EVE-0.5B 0.759 22.82 18.36 0.685 5.10 17.06
-EVE-7B 0.882 15.16 13.92 0.808 1.64 10.15
-Table 4: Performance comparison with two typical token merging methods.
-11 Visualization for Textual Region Compressor
-In the proposed EVE, the textual region compressor is introduced to perceive text-related information in each frame.
-To verify the effectiveness of this module, we have visualized the attention map of the textual region compressor in Fig.
-7.
-In the visualization, we observe that this module can indeed focus on the text region in each frame, which demonstrates the effectiveness of the proposed textual region compressor.
-Figure 7: The visualization for the textual region compressor.
-12 More Statistical Information of ViSa
-The detail information of SFT dataset have shown in the main text.
-In this section, we show some statistical information of the pre-training data in Fig.
-8.
-Most of pre-training videos contain less than 10 subtitles and their duration is between 11-20 seconds.
-13 More Result Visualizations
-We additionally visualize two comparison examples.
-Fig.9 and Fig.10 show the short video and movie examples, respectively.
-Through these two examples, we observe that Qwen2-VL and MiniCPM tend to describe the video content rather than predicting subtitles in SRT format.
-In addition, the cloud services can hardly extract bilingual subtitles, especially in movie videos.
-The ASR-based method PyVideoTrans also cannot achieve ideal performance since it only takes audio as input, which is not fully aligned with visual information.
-Compared with previous methods, the proposed EVE can better predict texts and timestamps of subtitles.
-Figure 8: The statistics information of ViSa’s pre-training dataset.
-Figure 9: A comparison example of short video.
-Figure 10: A comparison example of movie.
+Despite its strong performance, EVE still faces two limitations.
+First, although each frame is compressed into only 26 tokens, the maximum input length remains insufficient for very long videos, such as those exceeding one hour.
+Second, EVE predicts only subtitle texts and timestamps, without providing spatial coordinates of subtitle regions.
+Future work will explore more efficient token compression strategies to support longer input sequences and extend EVE to localize subtitle positions for broader downstream applications.